# Multi-scale Graph Autoregressive Modeling: Molecular Property Prediction via Next Token Prediction

### v1

@@ 4.2 Main Results
Table 2: Performance on MoleculeACE (RMSE). AVG: Average RMSE across all 30 tasks. Ranking: Bold (1st), Underline (2nd), Italic (3rd). Metric: $\downarrow$ denotes lower is better. Column: Previous Best Baselines denotes the best-performing method excluding KPGT; Type: ML (Machine Learning), DL (Deep Learning). Reference & Abbreviations: GMAE: GraphMAE (hou2022graphmae); MBRT: MoleBERT (xia2023mole); GROV: GROVER (rong2020self); SVM/RF/GBM: Support Vector Machine / Random Forest / Gradient Boosting Machine (van2022exposing). Descriptors: E: ECFP (rogers2010extended); M: MACCS (durant2002reoptimization).
@@ Appendix B Theoretical Derivations
XLNet points out two intrinsic issues of corruption-based MLM: (1) the special token corruption (e.g., [MASK]) creates a pretrain–finetune discrepancy, and (2) predicting multiple masked positions with a factorized objective effectively neglects their conditional dependencies given the unmasked context (yang2019xlnet). Relatedly, ELECTRA shows that much of the compute/sample-efficiency gap of MLM stems from defining the loss only on a small masked subset rather than all positions, while the [MASK] mismatch contributes a smaller but non-negligible part (clark2020electra).
@@
On graphs, the most predictive evidence for a token often lies in its local neighborhood (e.g., bond context, ring connectivity, functional-group surroundings). Random masking can therefore remove not only the target token but also its critical neighborhood evidence, making the missingness pattern itself a stochastic nuisance. This motivates a line of graph-pretraining designs that explicitly trade masking density for context stability: (1) masking structured units and adding global knowledge nodes (KPGT) (li2023knowledge), (2) masking local subgraphs and predicting contextual properties (GROVER) (rong2020self), and (iii) structure-guided / curriculum masking rather than purely random masking (StructMAE) (liu2024mask). These methods can be interpreted as efforts to find better operating points on the fundamental corruption–prediction trade-off of masked modeling (wettig2023should).
@@ B.2 Direct Supervision Density Analysis (Multi-Scale)
Increasing $\rho$ increases the number of predictions but also increases corruption, forming a fundamental trade-off in masked modeling (wettig2023should). In molecular graph Transformers, KPGT reports that setting the masking rate to $\rho=0.5$ can be optimal when token design and global knowledge tokens help stabilize the conditioning context (li2023knowledge). More broadly, structured masking / task design in graph pretraining (e.g., GROVER) can be viewed as attempts to obtain more useful target coverage without excessively destabilizing the evidence set (rong2020self).
@@
Dense token-level objectives are known to improve sample efficiency compared to losses defined only on a small masked subset; ELECTRA explicitly attributes a major part of MLM’s inefficiency to learning from only masked tokens rather than all positions (clark2020electra). CamS strengthens this further via multi-view repetition: the same molecule contributes token-level targets repeatedly across scales, inducing cross-scale consistency pressure and providing redundant evidence pathways (a motif may appear as fine tokens at small scales and as a merged token at larger scales). This reduces sensitivity to any single tokenization granularity and makes the effective supervision more robust.
@@ B.3 Structural Bias Formulation
The bias $\psi_{ij}$ is defined via static structural encodings (e.g., shortest-path distance), injected into attention as an additive bias term (ying2021transformers):
@@ Appendix D Detailed Experimental Results
Table 8: Full Performance comparison on MoleculeNet benchmark. (Extended version of Table 1). Sup variants: $_{\text{sup}}$ denotes adding an additional supervised graph-level bioactivity pre-training stage on top of the corresponding self-supervised objective (same protocol as used in the baseline setting). Baseline Reference: Infomax (velivckovic2018deep); Edgepred (hamilton2017inductive); Attribute Masking (Masking) (hu2020strategies); Context Prediction (Contextpred) (hu2020strategies); GraphLoG (xu2021self); GraphCL (you2020graph); JOAO (you2021graph); GROVER (rong2020self); 3DInfomax (stark20223d); GraphMVP (liu2016pyridazinone); ImageMol (zeng2022accurate); MolFormer (wu2023molformer); GEM (fang2022geometry); GraphMAE (hou2022graphmae); MoleBert (xia2023mole); KPGT (li2023knowledge)
@@
Table 9: Full Detailed results on 30 MoleculeACE activity-cliff tasks (Regression, RMSE $\downarrow$). (Extended version of Table 2). Sup variants: $_{\text{sup}}$ denotes adding an additional supervised graph-level bioactivity pre-training stage on top of the corresponding self-supervised objective (same protocol as used in the baseline setting). Descriptors: ECFP (rogers2010extended); MACCS (durant2002reoptimization); PHYSCHEM (walters2002prediction); WHIM (kubinyi19933d). Baseline Reference: support vector machines (SVM) (cristianini2002support); random forest (RF) (breiman1996bagging); gradient boosting machine (GBM) (friedman2001greedy); k-nearest neighbor (KNN) (fix1985discriminatory); message passing neural network (MPNN) (gilmer2017neural); graph attention network (GAT) (velivckovic2017graph); graph convolutional network (GCN) (kipf2016semi); AFP (xiong2019pushing); Convolutional Neural NetworkLong Short-Term Memory (CNN, specifically Maxsmi)  (kimber2021maxsmi); Long Short-Term Memory (LSTM, specifically CLM)  (moret2022perplexity); Transformer (specifically Chemberta) (chithrananda2020chemberta); KPGT (li2023knowledge)

### v2

@@ 4.2 Main Results
Table 2: Performance on MoleculeACE (RMSE). AVG: Average RMSE across all 30 tasks. Ranking: Bold (1st), Underline (2nd), Italic (3rd). Metric: $\downarrow$ denotes lower is better. Column: Previous Best Baselines denotes the best-performing method excluding KPGT; Type: ML (Machine Learning), DL (Deep Learning). Reference & Abbreviations: GMAE: GraphMAE (Hou et al., 2022); MBRT: MoleBERT (Xia et al., 2023); GROV: GROVER (Rong et al., 2020); SVM/RF/GBM: Support Vector Machine / Random Forest / Gradient Boosting Machine (Van Tilborg et al., 2022). Descriptors: E: ECFP (Rogers and Hahn, 2010); M: MACCS (Durant et al., 2002).
@@ Appendix B Theoretical Derivations
XLNet points out two intrinsic issues of corruption-based MLM: (1) the special token corruption (e.g., [MASK]) creates a pretrain–finetune discrepancy, and (2) predicting multiple masked positions with a factorized objective effectively neglects their conditional dependencies given the unmasked context (Yang et al., 2019). Relatedly, ELECTRA shows that much of the compute/sample-efficiency gap of MLM stems from defining the loss only on a small masked subset rather than all positions, while the [MASK] mismatch contributes a smaller but non-negligible part (Clark et al., 2020).
@@
On graphs, the most predictive evidence for a token often lies in its local neighborhood (e.g., bond context, ring connectivity, functional-group surroundings). Random masking can therefore remove not only the target token but also its critical neighborhood evidence, making the missingness pattern itself a stochastic nuisance. This motivates a line of graph-pretraining designs that explicitly trade masking density for context stability: (1) masking structured units and adding global knowledge nodes (KPGT) (Li et al., 2023), (2) masking local subgraphs and predicting contextual properties (GROVER) (Rong et al., 2020), and (iii) structure-guided / curriculum masking rather than purely random masking (StructMAE) (Liu et al., 2024). These methods can be interpreted as efforts to find better operating points on the fundamental corruption–prediction trade-off of masked modeling (Wettig et al., 2023).
@@ B.2 Direct Supervision Density Analysis (Multi-Scale)
Increasing $\rho$ increases the number of predictions but also increases corruption, forming a fundamental trade-off in masked modeling (Wettig et al., 2023). In molecular graph Transformers, KPGT reports that setting the masking rate to $\rho=0.5$ can be optimal when token design and global knowledge tokens help stabilize the conditioning context (Li et al., 2023). More broadly, structured masking / task design in graph pretraining (e.g., GROVER) can be viewed as attempts to obtain more useful target coverage without excessively destabilizing the evidence set (Rong et al., 2020).
@@
Dense token-level objectives are known to improve sample efficiency compared to losses defined only on a small masked subset; ELECTRA explicitly attributes a major part of MLM’s inefficiency to learning from only masked tokens rather than all positions (Clark et al., 2020). CamS strengthens this further via multi-view repetition: the same molecule contributes token-level targets repeatedly across scales, inducing cross-scale consistency pressure and providing redundant evidence pathways (a motif may appear as fine tokens at small scales and as a merged token at larger scales). This reduces sensitivity to any single tokenization granularity and makes the effective supervision more robust.
@@ B.3 Structural Bias Formulation
The bias $\psi_{ij}$ is defined via static structural encodings (e.g., shortest-path distance), injected into attention as an additive bias term (Ying et al., 2021):
@@ Appendix D Detailed Experimental Results
Table 8: Full Performance comparison on MoleculeNet benchmark. (Extended version of Table 1). Sup variants: $_{\text{sup}}$ denotes adding an additional supervised graph-level bioactivity pre-training stage on top of the corresponding self-supervised objective (same protocol as used in the baseline setting). Baseline Reference: Infomax (Veličković et al., 2018); Edgepred (Hamilton et al., 2017); Attribute Masking (Masking) (Hu et al., 2020); Context Prediction (Contextpred) (Hu et al., 2020); GraphLoG (Xu et al., 2021); GraphCL (You et al., 2020); JOAO (You et al., 2021); GROVER (Rong et al., 2020); 3DInfomax (Stärk et al., 2022); GraphMVP (Liu et al., 2016); ImageMol (Zeng et al., 2022); MolFormer (Wu et al., 2023); GEM (Fang et al., 2022); GraphMAE (Hou et al., 2022); MoleBert (Xia et al., 2023); KPGT (Li et al., 2023)
@@
Table 9: Full Detailed results on 30 MoleculeACE activity-cliff tasks (Regression, RMSE $\downarrow$). (Extended version of Table 2). Sup variants: $_{\text{sup}}$ denotes adding an additional supervised graph-level bioactivity pre-training stage on top of the corresponding self-supervised objective (same protocol as used in the baseline setting). Descriptors: ECFP (Rogers and Hahn, 2010); MACCS (Durant et al., 2002); PHYSCHEM (Walters and Murcko, 2002); WHIM (Kubinyi, 1993). Baseline Reference: support vector machines (SVM) (Cristianini and Scholkopf, 2002); random forest (RF) (Breiman, 1996); gradient boosting machine (GBM) (Friedman, 2001); k-nearest neighbor (KNN) (Fix, 1985); message passing neural network (MPNN) (Gilmer et al., 2017); graph attention network (GAT) (Veličković et al., 2017); graph convolutional network (GCN) (Kipf, 2016); AFP (Xiong et al., 2019); Convolutional Neural NetworkLong Short-Term Memory (CNN, specifically Maxsmi)  (Kimber et al., 2021); Long Short-Term Memory (LSTM, specifically CLM)  (Moret et al., 2022); Transformer (specifically Chemberta) (Chithrananda et al., 2020); KPGT (Li et al., 2023)